40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1116     Entry time: Thu Nov 6 09:45:27 2008
Author: steve 
Type: Update 
Category: MOPA 
Subject: head temp hick-up vs power  
The control room AC temp was lowered from 74F to 70F around Oct 10
This hold the head temp rock solid 18.45C for ~30 days as it shows on this 40 days plot.
We just had our first head temp hick-up

note: the laser chiller did not produce any water during this period
Attachment 1: htpr.jpg  67 kB  | Hide | Hide all
htpr.jpg
ELOG V3.1.3-